OpenStudy (anonymous):

8/(x^2-16)+3/4=1(x-4)

OpenStudy (anonymous):

\[ \frac{8}{x^2-16}+\frac 3 4=\frac 1 {x-4}\\ \frac{8}{(x-4)(x+4)}+\frac 3 4=\frac 1 {x-4}\\ \] Multiply the two sides by (x-4)(x+4)

OpenStudy (anonymous):

Multiply by 4(x-4)(x+4) \[ 32 + 3(x-4)(x+4) =4( x+4) \] Can you finish it from here?

OpenStudy (anonymous):

Expand both sides \[ 3 x^2-16=4 x+16\\ 3 x^2-4 x-32=0\\ (x-4) (3 x+8)=0\\ x=4\\ x=-\frac 83 \] x=4 is not acceptable/ Why?

OpenStudy (anonymous):

its extraneous and thank you for your help

OpenStudy (anonymous):

yw